Suppose that in the distant future, astronauts are exploring a planet in another solar system. They find that the radius of the planet is 6.8 x 103 km and the acceleration due to gravity on its surface is 15.3 m/s2. What is the mass of the planet?  


1
Expert's answer
2021-04-28T09:54:39-0400

By the law of universal gravitation:


"mg=\frac{GMm}{R^2},\\\space\\
M=\frac{gR^2}{G}=1.06·10^{25}\text{ kg}."